# Reinforcement Learning techniques for the flavor problem in particle physics

### v1

@@ 3.3 An AI agent for autonomous model building
As a step forward in response to the challenges posed by the flavor problem in the $\nu$ sector, the authors of [133] introduce the Autonomous Model Builder (AMBer), an AI-assisted framework based on RL. AMBer is designed to efficiently navigate the vast combinatorial landscape of model-building possibilities. In its implementation, the system autonomously selects symmetry groups, defines particle content, assigns group representations, and evaluates the resulting models. The goal is to construct viable models that both minimise the number of free parameters and reproduce experimental observations. While the framework has been demonstrated in the context of neutrino flavor mixing, the underlying methodology is general and can be adapted to other domains of theoretical physics.
@@
The RL algorithm chosen for AMBer is Proximal Policy Optimisation (PPO), which trains two neural networks: a policy network for selecting actions and a value network for estimating expected returns. The loss function includes value, policy, and entropy terms, balancing exploitation and exploration. Training is performed in parallel over 250 environments, each with a maximum episode length of 1000 steps. Evaluation of models occurs every few steps, with the evaluation frequency depending on the chosen symmetry group. The search is conducted in several theory spaces, such as $A_{4}\times Z_{4}$ with up to five flavons, $A_{4}\times Z_{N}$ with variable $N$, and $T_{19}\times Z_{4}$ with up to six flavons.
@@
Figure 2: Training variables of interest over time for searches in three spaces: $A_{4}\times\mathbb{Z}_{4}$ (top), $A_{4}\times\mathbb{Z}_{N}$ (middle), and $T_{19}\times\mathbb{Z}_{4}$ (bottom). Left column: evolution of $\chi^{2}$ in blue (where the curve indicates the median $\log_{10}{\chi^{2}}$ over all environments) and the mean number of parameters $\langle n_{p}\rangle$ as training progresses in orange. Right column: number of valid models in orange and good ($\chi^{2}\leq 10$ and $n_{p}\leq$ 7) models in blue. From [133].
@@ 4 Conclusions
In this short review, we have summarized the first applications of Reinforcement Learning (RL) to the flavor problem in particle physics. Traditional approaches to quark and lepton mass hierarchies face the difficulty of exploring a vast model space, which quickly becomes intractable with increasing symmetries and parameters. RL-based frameworks have demonstrated the ability to efficiently navigate these spaces, identifying viable Froggatt–Nielsen charge assignments for quarks and leptons, often rediscovering known models and, in some cases, uncovering new ones consistent with experimental data. In the lepton sector, RL results point toward a statistical preference for Normal Ordering of neutrino masses. More advanced frameworks, such as AMBer, further illustrate the potential of AI agents to autonomously construct Beyond Standard Model theories with minimal parameter dependence and predictive power. Although this field is still in its infancy, the promising results obtained so far indicate that RL may become a valuable tool in theoretical model building, complementing traditional approaches and opening new directions in the search for solutions to the flavor problem.

### v2

@@ 3.3 An AI agent for autonomous model building
As a step forward in response to the challenges posed by the flavor problem in the $\nu$ sector, the authors of [Baretz:2025zsv] introduce the Autonomous Model Builder (AMBer), an AI-assisted framework based on RL. AMBer is designed to efficiently navigate the vast combinatorial landscape of model-building possibilities. In its implementation, the system autonomously selects symmetry groups, defines particle content, assigns group representations, and evaluates the resulting models. The goal is to construct viable models that both minimise the number of free parameters and reproduce experimental observations. While the framework has been demonstrated in the context of neutrino flavor mixing, the underlying methodology is general and can be adapted to other domains of theoretical physics.
@@
The RL algorithm chosen for AMBer is Proximal Policy Optimisation (PPO) that learns a policy directly and updates it carefully so it does not change too much at once. It uses a value function to help judge how good actions are, learns only from data generated by its current behavior, and uses advantage estimates to make learning more stable. By reusing the same data for a few updates, PPO achieves reliable learning without becoming unstable. The loss function includes value, policy, and entropy terms, balancing exploitation and exploration. Training is performed in parallel over 250 environments, each with a maximum episode length of 1000 steps. Evaluation of models occurs every few steps, with the evaluation frequency depending on the chosen symmetry group. The search is conducted in several theory spaces, such as $A_{4}\times Z_{4}$ with up to five flavons, $A_{4}\times Z_{N}$ with variable $N$, and $T_{19}\times Z_{4}$ with up to six flavons.
@@
Figure 2: Training variables of interest over time for searches in three spaces: $A_{4}\times\mathbb{Z}_{4}$ (top), $A_{4}\times\mathbb{Z}_{N}$ (middle), and $T_{19}\times\mathbb{Z}_{4}$ (bottom). Left column: evolution of $\chi^{2}$ in blue (where the curve indicates the median $\log_{10}{\chi^{2}}$ over all environments) and the mean number of parameters $\langle n_{p}\rangle$ as training progresses in orange. Right column: number of valid models in orange and good ($\chi^{2}\leq 10$ and $n_{p}\leq$ 7) models in blue. From [Baretz:2025zsv].
@@ 4 Conclusions
In this short review, we have summarized the first applications of Reinforcement Learning (RL) to the flavor problem in particle physics. Traditional approaches to quark and lepton mass hierarchies face the difficulty of exploring a vast model space, which quickly becomes intractable with increasing symmetries and parameters. RL-based frameworks have demonstrated the ability to efficiently navigate these spaces, identifying viable Froggatt–Nielsen charge assignments for quarks and leptons, often rediscovering known models and, in some cases, uncovering new ones consistent with experimental data. In the lepton sector, RL results point toward a statistical preference for Normal Ordering of neutrino masses. More advanced frameworks, such as AMBer, further illustrate the potential of AI agents to autonomously construct Beyond Standard Model theories with minimal parameter dependence and predictive power. Although this field is still in its infancy, the promising results obtained so far indicate that RL may become a valuable tool in theoretical model building, complementing traditional approaches and opening new directions in the search for solutions to the flavor problem, as those based on modular symmetries and non-holomorphic modular forms.
In the future, we think that RL will have a large impact on the development of fully autonomous theory-design frameworks and that its ability in exploring more complex parameter spaces against simple random scan will be clarified.